OpenStudy (anonymous):

13. Suppose you work at a retail store that sells teddy bears. The amount of your weekly paycheck is a flat $200 for 40 hours of work plus $10 for each teddy bear you sell. Let x be the number of teddy bears that you sell and f(x) be the amount of money you earn each week. Please answer the following questions (5 points each): a. Write an expression for f(x) b. What is a reasonable domain for this function? c. What does f(10) mean in the context of this problem? d. How much money do you earn if you sell 25 teddy bears in a week? e. How many

OpenStudy (anonymous):

e. How many bears do you need to sell if you want to make $450 in one week?

OpenStudy (***[isuru]***):

hi, did u get through a , b, c, d, parts ?

OpenStudy (anonymous):

no i need all of em

OpenStudy (***[isuru]***):

Ok ... let's start with (a)

OpenStudy (anonymous):

ok

OpenStudy (***[isuru]***):

what is the amount u will get per a week except for the amount u get for the no, of bears u sell ?

OpenStudy (anonymous):

i thought it was f(x)=200+40+10x

OpenStudy (anonymous):

200 and 40 so 240 not counting the bears which is 10x

OpenStudy (***[isuru]***):

no!! the question says 200 for 40 hours it says that we will have to work 40 hours per week and we will get an amount of 200 as payback ... 40 mentioned here does not belong to the amount ...Did u get it ?

OpenStudy (anonymous):

yes

OpenStudy (***[isuru]***):

so.... the amount u will get for the work is 200 and u will get 10 dollars for each teddy bear u sell If u sell "x" amount of teddy bears in a week , u will get 10x amount of teddy bear bonus So the total amount u will earn a week is 200 + the amount of teddy bears u sell = 200 + 10x so the expression for the part (a) is f(x) = 200 + 10x got it ?

OpenStudy (anonymous):

yes

OpenStudy (***[isuru]***):

for the part (b) here f(x) = 200 + 10x can take any value for "x" which mean the domain of f(x) is the set Of REAL NUMBERS... but real numbers include fractions , negative and positive numbers ect. those thing doesn't exist in real world... I mean u can't find pieces of teddy bears or square root values of teddy bears or negative amount of teddy bears ect. So the range of values of teddy bears should only contain whole positive number like 0 , 1, 2 , 3, 4, .... So the reasonable domain for f(x) is the set of NATURAL numbers...

OpenStudy (***[isuru]***):

for the part (c)... f(10) means the total amount u get if u sell 10 teddy bears in a week

OpenStudy (***[isuru]***):

for the part (d). the total amount = 200 + no.of teddy bears u sell f(x) = 200 + 10x if u sell 25 teddy bears in a week f(25) = 200 + (25x10) = 200 + 250 = 450 so... u will earn $450 if u sell 25 teddy bears in a week

OpenStudy (***[isuru]***):

and the answer for part (d) is the question for part (e) (d) asks how much u will get if u sell 25 ; which is $450 (e) asks how many teddy bears u have to sell in order to get $450 it's25

OpenStudy (***[isuru]***):

Did ya get it ?

OpenStudy (anonymous):

yes thanks now e?

OpenStudy (***[isuru]***):

the question (e) and (d) r same.. u got 450 by selling 25 bears... and (e) asks how many to sell in order to get 450 ? u have to sell 25 ...

OpenStudy (anonymous):

thx

OpenStudy (***[isuru]***):

u r welcome!!!